1. Assume that the economic transactions described below occur during 201801, and reflect their implications on the NIPA tables below: - A lumber company produces and sells $200 of wood to a furniture producer. The furniture company produces a table with this wood, and sells it to a retailer for $500 - A household buys the table from the retail store for $600, and use it at home. Expenditure Account (Hint: include only final goods!) Consumption Investment Government expenditure Net Exports Gross Domestic Product Value Added by Sector Lumber (forestry) Furniture (manufacturing) Retail trade Gross Domestic Product Suppose, at each stage of production, half of the value added was paid to labor as wages and salaries, and the rest accrued as profit to the corporations. Fill in the income account below. Income Account Wages and salaries (labor) Corporate profits (capital) Gross Domestic Income

Need help, please show work
Answer #1

Expenditure Account

Consumption = 600 ( the final value of good consumed by the household. All the other products were not consumed. Rather they were being processed for final use)

Investment = 0 (There is no investment made. An example of investment would have been lumber buying a new machine)

Government Expenditure = 0. (There has been no govt. expenditure)

Net Exports =0. ( There is no exports or imports)

Hence, GDP = 600 + 0 + 0 + 0 = 600

Value added by sector

Lumber - 200 (he produced wood worth $200)

Manufacturer - 300 (he converted wood worth $200 into furniture worth $500)

Retail trade - 100 (the value of services given by retailer = 600-500)

Hence, GDP = 200 + 300 + 100 = 600

Income Method

Wages and salaries = 100 + 150 + 50 = 300 ( half of value added at each stage)

Corp. Profit = 100 + 150 + 50 = 300 (half of value added at each stage)

Hence, GDP = 300 + 300 = 600